The condensed form of Nordstrop Company’s flexible budget for manufacturing overhead follows:

Cost Machine-Hours
Formula
Overhead Costs (per machine-hour) 7,500 8,500 9,500
  Variable cost $ 1.35 $ 10,125   $ 11,475   $ 12,825  
  Fixed cost 15,000   15,000   15,000  
  Total overhead cost $ 25,125   $ 26,475   $ 27,825  

The following information is available for a recent period:

  1. The denominator activity of 7,500 machine-hours was chosen to compute the predetermined overhead rate.
  2. At the 7,500 standard machine-hours level of activity, the company should produce 3,750 units of product.
  3. The company’s actual operating results were as follows:
  
  Number of units produced 4,050  
  Actual machine-hours 8,200  
  Actual variable overhead costs $ 12,000  
  Actual fixed overhead costs $ 16,000  

Required:

1. Compute the predetermined overhead rate and break it down into variable and fixed cost elements. (Round your answers to 2 decimal places.)

2. What were the standard hours allowed for the year’s actual output? (Do not round intermediate calculations.)

3. Compute the variable overhead spending and efficiency variances and the fixed overhead budget and volume variances. (Do not round intermediate calculations. Round your answers to 2 decimal places. Indicate the effect of each variance by selecting "F" for favourable, "U" for unfavourable, and "None" for no effect (i.e., zero variance).)
Answer #1

Answers

  • Requirement 1

>Predetermined Overhead rate = $ 1.35 + ($15000 / 7500 machine hours) = 1.35 + 2 = $ 3.35 per machine hours
>variable cost rate = $ 1.35 per machine hour
>fixed cost rate = $ 2.00 per machine hour

  • Requirement 2
    Standard hours allowed = 4050 units x (7500 hours / 3750 units) = 8100 machine hours
  • Requirement 3

Variable Overhead Spending Variance

(

Standard Cost

-

Actual Cost

)

(

$                      10,935.00

-

$            12,000.00

)

-1065

Variance

$              1,065.00

Unfavourable-U

Variable Overhead Efficiency Variance

(

Standard Hours

-

Actual Hours

)

x

Standard Rate

(

8100

-

8200

)

x

$                           1.35

-135

Variance

$                  135.00

Unfavourable-U

Fixed Overhead Production Budget Variance

(

Budgeted Fixed Overhead

-

Actual Fixed Overhead incurred

)

(

$                      15,000.00

-

$            16,000.00

)

-1000

Variance

$              1,000.00

Unfavourable-U

Fixed Overhead Production Volume Variance

(

Standard Fixed Overhead or Fixed Overhead absorbed

-

Budgeted Fixed Overhead

)

(

$                      16,200.00

-

$            15,000.00

)

1200

Variance

$              1,200.00

Favourable-F